Find the spherical coordinates of the point whose coordinates are
1) (1, 1, √6)
2) (-2, 2√3, 4)
3) (-√3, 1, -2)
4) (4,-4√3, 6)
1
Expert's answer
2020-12-17T18:40:12-0500

We remind the expressions of spherical coordinates :

"\rho = \sqrt{x^2+y^2+z^2}"

"\theta = \arccos(\frac{z}{\rho})"

"\phi" such that "x=\rho \sin(\theta) \cos(\phi), y=\rho\sin(\theta)\sin(\phi)"

Therefore we have :

  1. "(\rho,\theta,\phi) = (2\sqrt{2}, \frac{\pi}{6}, \frac{\pi}{4})"
  2. "(4\sqrt{2}, \frac{\pi}{4}, \frac{2\pi}{3})"
  3. "(2\sqrt{2}, \frac{3\pi}{4}, \frac{5\pi}{6})"
  4. "(10, \arccos(\frac{3}{5}), -\frac{\pi}{3})"